# Controllable non-Hermitian qubit-qubit Coupling in Superconducting quantum Circuit

### v1

@@ III Controllable Non-Hermitian coupling
The larger oscillating amplitudes in the black-dash-dotted curves for $\Delta\theta\approx\pi/2$ originates from the small damping rate contributed by non-hermite coupling close to exceptional points. The maximal amplitude of non-hermite coupling strength is defined as $\Omega_{n}=\lambda_{1}\lambda_{2}/\gamma_{a}$, the positions of local maximal or minimal values for the envelope are different on the four curves which originate from the dependencies of total qubit-qubit coupling on the J-C coupling phases. The different values of local maximal values on envelope imply that the efficient of quantum state exchange between two qubits can be tuned by the J-C coupling phase, and the exchange efficient gets the highest value close to the exception point. In the black curves ($\Delta\theta=\pi/2$), the envelope of occupation probability drops slowly compared with the results on the blue-solid ($\Delta\theta=0$), green-dashed ($\Delta\theta=\pi/3$), and red dash-dotted ($\Delta\theta=5\pi/12$) curves, which indicates the different contributions of the non-hermite coupling on the effective damping rates of qubits. The quantum state exchange efficient and the costing time getting local maximal values can be tuned with the J-C coupling phases which might be useful to enhance the speed and fidelity of two-qubit quantum gates.
IV nonreciprocal coupling in Non-Hermitian circuit
@@
As indicted by the previous work, the non-reciprocal transmission can be realized in non-Hermitian Circuit[46, 47, 48]. According to Eq.(3), the non-Hermitian coupling strengths between two qubits in the forward and backward directions can be respectively defined as $g^{(q)}_{1\rightleftharpoons 2}=g_{e}-i\Omega_{n}\exp{(\mp i\Delta\theta)}$, where the first term $g_{e}$ is hermite interactions, while the second-term $i\Omega_{n}\exp{(\mp i\Delta\theta)}$ is the non-hermite qubit-qubit coupling induced by the high-loss resonator.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
Figure 4: (Color online) Controllable Non-reciprocal couplings and state evolution in non-Hermitian superconducting circuit. (a) Non-Hermitian coupling strengths $g^{(q)}_{1\rightarrow 2}$ and (b) $g^{(q)}_{1\leftarrow 2}$ as functions of phase difference ($\Delta\theta$) and effective coupling strength ($g_{e}$). The one-dimensional curves of $g^{(q)}_{1\rightarrow 2}$ (blue-dashed curve) and $g^{(q)}_{1\leftarrow 2}$ (red-solid curve) for (c) $\Delta\theta=\pi/3$ and (e) $\Delta\theta=\pi/2$. The one-dimensional curves of non-reciprocal ratio $|g^{(q)}_{1\leftarrow 2}|/|g^{(q)}_{1\rightarrow 2}|$ for (d) $\Delta\theta=\pi/2$ and (f) $\Delta\theta=\pi/3$. The two qubits are assumed as in the ground states, $\langle\sigma^{(j)}_{z}\rangle=-1$ with $j=1,2$. The other parameters in (a)-(d) are: $\omega_{a}/2\pi=4.475$ GHz, $\omega_{1}/2\pi=4.5$ GHz, $\omega_{2}/2\pi=4.505$ GHz, $\omega^{(max)}_{c}/2\pi=5.20$ GHz, $\gamma_{1}/2\pi=1.00$ MHz, $\gamma_{2}/2\pi=1.01$ MHz, $g_{xy}/2\pi=4.0$ MHz, $\lambda_{j}/2\pi=11$ MHz, $g_{1}/2\pi=30$ MHz, $g_{2}/2\pi=30.3$ MHz, $g_{e}/2\pi=1.137$ MHz, and $\gamma_{a}/2\pi=65$ MHz. The differences in occupation probabilities for two qubits starting in (g) excited states ($P^{(e)}_{j}$) and (h) ground ($P^{(g)}_{j}$) states. The two qubits adopt same parameters:$\omega_{j}/2\pi=4.5$ GHz, $g_{j}/2\pi=30$ MHz, and $\gamma_{j}/2\pi=1$ MHz (with $j=1,2$), and the other parameters are the same as in (a)-(f) except for $g_{e}/2\pi\approx-80$ KHz.
@@
The $g^{(q)}_{1\rightarrow 2}\neq g^{(q)}_{1\leftarrow 2}$ even in the case of $\Delta\theta=n\pi$ and the local maximal (or minimal) values of $|g^{(q)}_{1\rightarrow 2}|$ and $|g^{(q)}_{1\leftarrow 2}|$ are out of steps as shown in Figs. 4(a) and 4(b), which implies the non-reciprocal coupling for two different directions of effective qubit-qubit coupling. The minimal values of $|g^{(q)}_{1\rightarrow 2}|$ and $|g^{(q)}_{1\leftarrow 2}|$ could reach zero in Fig. 4(e) for $\Delta\theta=\pi/2$, but they get nonzero minimal value in the case of $\Delta\theta=\pi/3$ in Fig. 4(c). And the nonreciprocal ratio is much stronger in the case of $\Delta\theta=\pi/2$ as indicated by the maximal and minimal values of $|g_{1\leftarrow 2}|/|g_{1\rightarrow 2}|$ in Figs. 4(d) and  4(f). The non-reciprocal coupling strengths between two qubits can also be tuned by the J-C coupling amplitudes and qubits’ quantum states.
@@
The nonreciprocal coupling between two qubits might also result in the asymmetry of quantum state evolutions from different directions, which can be seen from the nonzero values for difference in occupation probabilities of two qubits’ excited states (same initial states) in Figs. 4(g) and  4(h). Beside the apparently asymmetric of occupation probabilities after the period $t>1/\Omega_{n}$, the nonreciprocity of qubit quantum state evolution can also be tuned by the J-C coupling phases. Even adopting the same frequencies, same coupling strengths, same damping rates, and same initial states, the time evolution of qubit occupation probabilities from two different directions still do not overlap in Figs. 4(g) and  4(h), which coincides with the results of non-reciprocal coupling in Figs. 4(c) and  4(f).
@@ V Conclusions
In conclusion, the energy level degeneracies and quantum state evolutions are studied in non-Hermitian superconducting circuit. The non-hermite and nonreciprocal coupling between two qubits also affect the evolution of qubits’ quantum states, the efficient and asymmetry of quantum state exchange can be tuned by the J-C coupling phase which might be useful for the quantum gates. The resonator and qubits are in weak dispersive coupling regimes, so the state leakages and crosstalk are inevitable which should be suppressed with tunable coupler and wave-shape correction techniques[42].

### v2

@@ III Controllable Non-Hermitian coupling
The larger oscillating amplitudes in the black-dash-dotted curves for $\Delta\theta\approx\pi/2$ originates from the small damping rate contributed by non-hermite coupling close to exceptional points. The maximal amplitude of non-hermite coupling strength is defined as $\Omega_{n}=\lambda_{1}\lambda_{2}/\gamma_{a}$, the positions of local maximal or minimal values for the envelope are different on the four curves which originate from the dependencies of total qubit-qubit coupling on the phase difference $\Delta\theta$. The different values of local maximal values on envelope imply that the efficient of quantum state exchange between two qubits can be tuned by the resonator-qubit coupling phase, and the exchange efficient gets the highest value close to the exception point. In the black curves ($\Delta\theta=\pi/2$), the envelope of occupation probability drops slowly compared with the results on the blue-solid ($\Delta\theta=0$), green-dashed ($\Delta\theta=\pi/3$), and red dash-dotted ($\Delta\theta=5\pi/12$) curves, which indicates the different contributions of the non-hermite coupling on the effective damping rates of qubits. The quantum state exchange efficient and the costing time getting local maximal values can be tuned with the resonator-qubit coupling phases which might be useful to enhance the speed and fidelity of two-qubit quantum gates.
IV nonreciprocal coupling
@@
The non-reciprocal coupling and photon transmission can also be realized in non-Hermitian system[44, 45, 46, 47, 50, 51]. According to Eq.(3), the non-Hermitian coupling strengths between two qubits in the forward and backward directions can be respectively defined as $g^{(q)}_{1\rightleftharpoons 2}=g_{e}-i\Omega_{n}\exp{(\mp i\Delta\theta)}$, and the first and second term label the Hermite and non-Hermite type qubit-qubit coupling, respectively.
@@
Figure 4: (Color online) Controllable non-reciprocal couplings and state evolutions in non-Hermitian superconducting circuit. (a) Non-Hermitian coupling strengths $g^{(q)}_{1\rightarrow 2}$ and (b) $g^{(q)}_{1\leftarrow 2}$ as functions of phase difference $\Delta\theta$ and coupling strength $g_{e}$. The one-dimensional curves of $g^{(q)}_{1\rightarrow 2}$ (blue-dashed curve) and $g^{(q)}_{1\leftarrow 2}$ (red-solid curve) for (c) $\Delta\theta=\pi/3$ and (e) $\Delta\theta=\pi/2$. The non-reciprocal ratio $|g^{(q)}_{1\leftarrow 2}|/|g^{(q)}_{1\rightarrow 2}|$ for (d) $\Delta\theta=\pi/2$ and (f) $\Delta\theta=\pi/3$. The two qubits are assumed as in the ground states, $\langle\sigma^{(j)}_{z}\rangle=-1$ with $j=1,2$. The other parameters in (a)-(d) are: $\omega_{a}/2\pi=4.475$ GHz, $\omega_{1}/2\pi=4.5$ GHz, $\omega_{2}/2\pi=4.505$ GHz, $\omega^{(max)}_{c}/2\pi=5.20$ GHz, $\gamma_{1}/2\pi=1.00$ MHz, $\gamma_{2}/2\pi=1.01$ MHz, $g_{xy}/2\pi=4.0$ MHz, $\lambda_{j}/2\pi=11$ MHz, $g_{1}/2\pi=30$ MHz, $g_{2}/2\pi=30.3$ MHz, $g_{e}/2\pi=1.137$ MHz, and $\gamma_{a}/2\pi=65$ MHz. The differences in occupation probabilities of two qubits starting in (g) excited states ($P^{(e)}_{j}$) and (h) ground ($P^{(g)}_{j}$) states. $\omega_{j}/2\pi=4.5$ GHz, $g_{j}/2\pi=30$ MHz, and $\gamma_{j}/2\pi=1$ MHz, with $j=1,2$,, and the other parameters are the same as in (a)-(f) except for $g_{e}/2\pi\approx-80$kHz.
Figure 4: (Color online) Controllable non-reciprocal couplings and state evolutions in non-Hermitian superconducting circuit. (a) Non-Hermitian coupling strengths $g^{(q)}_{1\rightarrow 2}$ and (b) $g^{(q)}_{1\leftarrow 2}$ as functions of phase difference $\Delta\theta$ and coupling strength $g_{e}$. The one-dimensional curves of $g^{(q)}_{1\rightarrow 2}$ (blue-dashed curve) and $g^{(q)}_{1\leftarrow 2}$ (red-solid curve) for (c) $\Delta\theta=\pi/3$ and (e) $\Delta\theta=\pi/2$. The non-reciprocal ratio $|g^{(q)}_{1\leftarrow 2}|/|g^{(q)}_{1\rightarrow 2}|$ for (d) $\Delta\theta=\pi/2$ and (f) $\Delta\theta=\pi/3$. The two qubits are assumed as in the ground states, $\langle\sigma^{(j)}_{z}\rangle=-1$ with $j=1,2$. The other parameters in (a)-(d) are: $\omega_{a}/2\pi=4.475$ GHz, $\omega_{1}/2\pi=4.5$ GHz, $\omega_{2}/2\pi=4.505$ GHz, $\omega^{(max)}_{c}/2\pi=5.20$ GHz, $\gamma_{1}/2\pi=1.00$ MHz, $\gamma_{2}/2\pi=1.01$ MHz, $g_{xy}/2\pi=4.0$ MHz, $\lambda_{j}/2\pi=11$ MHz, $g_{1}/2\pi=30$ MHz, $g_{2}/2\pi=30.3$ MHz, $g_{e}/2\pi=1.137$ MHz, and $\gamma_{a}/2\pi=65$ MHz. The differences in occupation probabilities of two qubits starting in (g) excited states ($P^{(e)}_{j}$) and (h) ground ($P^{(g)}_{j}$) states. $\omega_{j}/2\pi=4.5$ GHz, $g_{j}/2\pi=30$ MHz, and $\gamma_{j}/2\pi=1$ MHz, with $j=1,2$,, and the other parameters are the same as in (a)-(f) except for $g_{e}/2\pi\approx-80$kHz.
@@
The $g^{(q)}_{1\rightarrow 2}\neq g^{(q)}_{1\leftarrow 2}$ even in the case of $\Delta\theta=n\pi$ and the local maximal (or minimal) values of $|g^{(q)}_{1\rightarrow 2}|$ and $|g^{(q)}_{1\leftarrow 2}|$ are out of steps as shown in Figs. 4(a) and 4(b), which implies the non-reciprocal coupling for two different directions of effective qubit-qubit coupling. The minimal values of $|g^{(q)}_{1\rightarrow 2}|$ and $|g^{(q)}_{1\leftarrow 2}|$ could reach zero in Fig. 4(e) for $\Delta\theta=\pi/2$, but they get nonzero minimal value in the case of $\Delta\theta=\pi/3$ in Fig. 4(c). And the nonreciprocal ratio is much stronger in the case of $\Delta\theta=\pi/2$ as indicated by the maximal and minimal values of $|g_{1\leftarrow 2}|/|g_{1\rightarrow 2}|$ in Figs. 4(d) and  4(f). The non-reciprocal coupling strengths between two qubits can also be tuned by the resonator-qubit coupling amplitudes and qubits’ quantum states.
@@
The nonreciprocal coupling between two qubits might result in the asymmetry of quantum state evolutions for different directions, which can be seen from the nonzero values of occupation probability difference for two qubits’ excited states with the same initial states in Figs. 4(g) and  4(h). Beside the apparently asymmetric of occupation probabilities after the period $t>1/\Omega_{n}$, the nonreciprocity of qubit quantum state evolution can also be tuned by the resonator-qubit coupling phases. Even adopting the same frequencies, same coupling strengths, same damping rates, and same initial states, the time evolution of qubit occupation probabilities from two different directions still do not overlap in Figs. 4(g) and  4(h), which coincides with the results of non-reciprocal coupling in Figs. 4(c) and  4(f).
@@ V Conclusions
In conclusion, the energy level degeneracies and quantum state evolutions are studied in non-Hermitian superconducting circuit. The efficient and asymmetry of quantum state evolutions can be tuned by the resonator-qubit coupling phase which might be useful for enhancing the fidelity and speeding of quantum gates. The resonator and qubits are in weak dispersive coupling regimes, so the state leakages and crosstalks are inevitable which should be suppressed with tunable coupler and wave-shape correction techniques[40].